Opened 9 years ago

Closed 7 years ago

Last modified 7 years ago

#355 closed flight processing (fixed)

EUFAR10/03 (A.NEW), flight day 197/2010,Lisbon

Reported by: ella Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: emca

Description (last modified by iopa)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EUFAR10_03-2010_197_Lisbon

Data arrived from ARSF via network transfer on 2nd/08/2010.

Scientific objectives: Objective 1.- Development of a reliable tool to estimate coastal water quality indicators by using remote sensing
techniques, both airborne and satellite.
1.1. | Development of a set of semi-empirical algorithms based on hyper-spectral airborne and in-situ data for
estimating the concentrations of the water constituents;
1.2. | Migration of the developed algorithms into the satellite workflow.

Priority: a4h

PI: Jose da Silva

Sensors:

  • Eagle (09/03/2011)
  • Hawk (09/03/2011)
  • Leica LIDAR (03/12/2010)
  • RCD Camera (03/12/2010)

Attachments (2)

line_split_times (513 bytes) - added by iopa 8 years ago.
Eagle & Hawk GPS Time splits
1972010.sync (228 bytes) - added by iopa 8 years ago.
Eagle & Hawk SCT offsets

Download all attachments as: .zip

Change History (43)

comment:1 Changed 9 years ago by ella

  • Description modified (diff)

comment:2 Changed 9 years ago by ella

  • Description modified (diff)

comment:3 Changed 9 years ago by ella

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:4 Changed 9 years ago by jb362

Starting hyperspectral processing

comment:5 Changed 9 years ago by jb362

Leaving this hyperspectral processing to be picked up again once we have a logsheet for the flight. Applanix navigation has been done.

comment:6 Changed 9 years ago by jb362

Starting camera processing.

comment:7 Changed 9 years ago by ella

IPAS navigation now processed.

comment:8 Changed 9 years ago by jb362

Trying hyperspectral processing again without the logsheet.

comment:9 Changed 9 years ago by jb362

Doesn't look like this will work so will have to wait for logsheet still.

comment:10 Changed 9 years ago by jb362

Ignore last comment!

comment:11 Changed 9 years ago by jb362

The eagle lev3 flight lines for this project do not appear to match up with the kml file when opened in gtviewer. We have decided to wait for the logsheets to see if those can shed any more light on the situation. Also, this flight is almost entirely over the ocean (with only a tiny piece of land visible on one flight line) and therefore (1) will not require a DEM and (2) will be impossible to find SCT offsets for.

comment:12 Changed 9 years ago by jb362

Photographs 1-22 removed as they did not cover the study area.

comment:13 Changed 9 years ago by emca

Logsheet has arrived and been saved in ~arsf/arsf_data/2010/flight_data/archive-overflow/EUFAR10_03-2010_197_Lisbon/admin

comment:14 Changed 9 years ago by jb362

Iopa will be taking over delivery of this for me.

comment:15 Changed 9 years ago by ella

RCD delivery check complete, ready to deliver.

comment:16 Changed 9 years ago by iopa

  • Description modified (diff)

RCD camera was dispatched today, 16th September 2010.

comment:17 Changed 8 years ago by knpa

  • Description modified (diff)

Moved to repository, can be rsynced.

comment:18 Changed 8 years ago by benj

  • Description modified (diff)

Corrected PI (recorded wrong on data unpacking)

comment:19 Changed 8 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:20 Changed 8 years ago by iopa

Taking eagle.

comment:21 Changed 8 years ago by emca

  • Other processors set to emca

Starting LiDAR processing

comment:22 Changed 8 years ago by emca

LiDAR processing complete. Ready for delivery check at
~airborne/workspace/EUFAR10_03-2010_197_Lisbon/delivery/20101112/EUFAR10_03

LiDAR data was collected as one continuous line which has been split to match approximately with the Eagle/Hawk data. It was not possible to correct for roll and no DEM was created for the delivery as this project is entirely over water.

comment:23 Changed 8 years ago by benj

  • Description modified (diff)

comment:24 Changed 8 years ago by iopa

LiDAR has been checked,

no errors in the data or the delivery, ready to go.

comment:25 Changed 8 years ago by emca

  • Description modified (diff)

LiDAR and Photography dispatched 03/12/2010

(note: photography is a re-delivery as data sent on 16/09/10 was sent to the wrong PI)

comment:26 Changed 8 years ago by iopa

Taking Hawk processing for this as well.

comment:27 Changed 8 years ago by iopa

Hyperspectral status:

The flightlines for both Eagle and Hawk have contained very large amounts of turns and to process them correctly they have been split into parts to exclude the major >75 degree turns. The precise splitting points are given in the file line_split_times which contains GPS times that apply to both Eagle and Hawk.

Because of this however, some parts had to be processed with has_sync = false (excluding the parts whose start was the same as their original line's start) and so due to the approximation of their header times there will be some timing errors, which also cannot be corrected to the usual precision since the flight is fully over water.

comment:28 Changed 8 years ago by iopa

To utilise those GPS times the command splitline.py was used.

Changed 8 years ago by iopa

Eagle & Hawk GPS Time splits

Changed 8 years ago by iopa

Eagle & Hawk SCT offsets

comment:29 Changed 8 years ago by iopa

Hyperspectral is ready for delivery check.

While both Eagle and Hawk have been split into parts to be processed with azgcorr, no offsets were applied on Eagle as there are no reference points to de-wobble. Hawk was offset appropriately to match the Eagle, using sun glints, clouds or other odd bits.

comment:30 Changed 8 years ago by knpa

  • Owner set to iopa
  • Status changed from new to assigned

Taking HS delivery check.

comment:31 Changed 8 years ago by knpa

Delivery check complete.

All looks good apart from the following:

  1. DEM directory is empty. If we have one then put it in here, otherwise remove dir and reference in Readme.
  2. Something wrong with this sentence in Quality issues section: "During the production of the data, we processed Level 3 imagery . Screenshots of the Level 3"
  3. azgcorr command - should have an example of dem even though we don't provide one. Unless it's all over water then obviously they don't need a dem - in which case I don't think we need the -es command either as no geoid-spheroid sep needs to be done? Maybe shouldn't include the misc directory either.
  4. Couldn't open mosaic screenshots in eog (too big). Is there a way to reduce their size?
  5. Regular dimming and brightening of Hawk throughout all flightlines. Should mention in Readme (and maybe analyse - seen this quite a lot but never heard it mentioned before).
  6. Underflow and overflow on parts of all eagle lines, caused by sunglint. Huge underflows on Hawk caused by the above effect. Mention in readme.

comment:32 Changed 8 years ago by iopa

Delivery check comments have been amended, a reduced size version for each mosaic has been produced and included as well as the original. More quality remarks have been added.

Delivery has been put on hold to investigate point no 5 of Hawk dimming/brightening.

comment:33 Changed 8 years ago by iopa

  • Description modified (diff)

Hyperspectral has been sent to PI on 9th March 2011.

comment:34 Changed 8 years ago by anch

  • Summary changed from EUFAR10/03, flight day 197/2010,Lisbon to EUFAR10/03 (A.NEW), flight day 197/2010,Lisbon

comment:35 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from iopa to knpa

rsynced project.

Beginning archiving.

comment:36 Changed 8 years ago by knpa

Tabi data present.

comment:37 Changed 8 years ago by knpa

Tarball created. Made available to NEODC.

comment:38 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion

Backed up by NEODC. Moved to archive disk (this happened a while ago, forgot to update ticket).

comment:39 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:40 Changed 7 years ago by knpa

Delivery checking APL reprocessing of HS.

comment:41 Changed 7 years ago by knpa

Delivery check complete.

All looks good apart from:

  • gdal_translate example command in ReadMe has got munged (pg 4)
Note: See TracTickets for help on using tickets.